a) A patient is found to have a urine flow rate of 1 ml/min, a plasma urea concentration of 20 mg/100 ml, and a urinary urea concentration of 1000 mg/100 ml. Calculate the specific plasma clearance of urea for this individual.

b)What is the relative plasma clearance of urea for the patient in “A” if GFR = 75 ml/min? Does the patient reabsorb or secrete urea? How can you tell?
Answer #1

a) specific plasma clearance of urea: cm= U×V/P

Cm- Maximum urea clearance.

U- Urea concentration in urine = 1000mg/100ml,

that is, 10mg/1ml

V- Urea excreted per minute per ml. 75ml/minute

P- Urine concentration in plasma. 20mg/100ml

That is, 0.2mg/ ml

Cm= 10×75/0.2

= 750/0.2

= 32.5

Maximum area clearance 32.5.

B) Normal GFR rate= more than 90ml / minutes

If its 75ml/ minutes indicates plasma clearance of patient is reduced.

Urea reabsorption is reduced.

Its indicated by GFR rate